You manage a real estate investment company. One year ago, the company purchased 10 parcels of land distributed throughout the community for $ 12.0 million each. A recent appraisal of the properties indicates that five of the parcels are now worth $9.8 million each, while the other five are worth $18.0 million each. Ignoring any income received from the properties and any taxes paid over the year, calculate the investment company’s accounting earnings and its economic earnings in each of the following cases:

The company sells all of the properties at their appraised values today.

The company sells none of the properties.

The company sells the properties that have fallen in value and keeps the others.

The company sells the properties that have risen in value and keeps the others.
Answer #1

Economic Income is the income which is increase over time over its initial value wether it is sold or not. Economic Income is same in all cases since it is not the realized income but the income in economic term due to apprised value.

Economic Income = 9.8*5 + 18*5 - 12*10

= $19

Accounting income is the income equla to the selling price minus cost price.

Case A) Accounting income = 9.8*5 + 18*5 - 12*10

= $19

Case B) Since no transaction

Accounting Income = $0

Case C) Accounting income = 9.8*5 - 12*5

= -$11

Case D) Accounting Income = 18 * 5 - 12*5

= $30

Case Accounting Income Economic Income
1 $ 19.00 $ 19.00
2 $                           -   $ 19.00
3 $                   (11.00) $ 19.00
4 $                     30.00 $ 19.00